Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 13

1. Section 15.1: Problems 9, 15, 17, 25, 27, 28.

2. Section 15.2: Problems 5, 17. (‘Work’, ‘Flow’, ‘Circulation’ are alternative names for the integral ∫ F · T ds, see Definition at p939 and p941. Also read Definition at p942 and then p943-944 for the definition and formula for ‘flux’ for future use).